OpenStudy (anonymous):

find the inverse of the function defined by f (x) = x+5/x-2

OpenStudy (anonymous):

\[y=\frac{x+5}{x-2}\] Switching variable \[x=\frac{y+5}{y-2}\] \[x(y-2)=y+5\] \[xy-2x=y+5\] \[xy-y=5+2x\] \[y(x-1)=5+2x\] \[y=\frac{5+2x}{(x-1)}\]